A firm in a monopoly market is facing a demand curve P=100-2x , total revenue curve TR=10x2x2 and has total cost C=10+2x.
a) Find the profit maximizing output and price?
b) Is this firm making a profit? Calculate the profit.
c) Show this on a graph.
1
Expert's answer
2020-06-11T11:19:38-0400

a)For a monopoly profit maximization

"MR=MC"

"P=100-2x"

"TR=P*Q"

"=(100-2x)x"

"=100x-2x^2"

"MR=" "dTR \over dQ" "=100-4x"

"TC=10+2x"

"MC=" "dTC \over dQ" "=2"

"MR=MC"

"100-4x=2"

"4x=98"

"x=24.5(output)"

Price

"P=100-2x"

"100-2(24.5)=51"

b) Yes

Profit "=TR-TC"

"TR=100(24.5)-2(24.5)^2"

"=1249.50"

"TC=10+2x"

"10+2(24.5)=59"

"1249.50-59=1190.5"